Which equation describes a line with a slope of 7 and includes the point ( −45, 65)?

Answer by lwsshak3(11628)

Which equation describes a line with a slope of 7 and includes the point ( −45, 65)?
***
use point-slope formula:y-y1=m(x-x1)
y-65=7(x-(-45))
y-65=7(x+45)
y-65=7x+315
y=7x+380